-- GiovanniCresci - 11 Dec 2009

Observers: Manolis, Giovanni, Dave, Nancy, Walter, Marco, Olga

00:20 UT --> sky partly cloudy at sunset, with clouds moving towards Mt. Graham.

01:20 UT --> sky cloudy, waiting for clear sky

03:20 UT --> opening the dome

03:56 UT --> IC5298 23:16:00.89 +25:33:29.4 Imaging, Bry, seeing 0.9", 1350s, thin cirrus

04:28 UT --> IC5298 23:16:00.89 +25:33:29.4 Imaging, H2, seeing 0.8", 1350s, thin cirrus

05:25 UT --> LBT_002 03:02:12.0 -00:01:33 Imaging, K, seeing 0.85", 1800s, very thin cirrus

06:15 UT --> high humidity, dome closed

06:40 UT --> opening the dome

06:45 UT --> SBS0335-052E 03:37:44 -05:02:30 Longslit, 1.93/200, seeing 0.8", aborted during acquisition

06:55 UT --> LUCIFER problem with flexure compensation

07:00 UT --> Seeing went bad (1.6") after acquisition, aborted

07:10 UT --> S255NS1 06:12:53.49 +18:00:11.6 Longslit, 2.20/210. seeing 1.2", clear

07:20 UT --> LUCIFER problem with flexure compensation, aborted

08:06 UT --> S255NS1 06:12:53.49 +18:00:11.6 Longslit, 2.20/210. seeing 1.3", 720s, clear

08:38 UT --> flat, arcs, telluric

09:20 UT --> J0901+1814 09:01:22.62 +18:14:24.6 Longslit,2.25/210, seeing 0.8" , clear

10:00 UT --> high humidity, dome closed after a frame affected by condensation

10:44 UT --> opening the dome again

10:50 UT --> J0901+1814 09:01:22.62 +18:14:24.6 Longslit,2.25/210, seeing 0.75", cirrus to the east

11:15 UT --> communication problem with the guide probe after 2 frames, no guiding, aborted

11:45 UT --> J0901+1814 09:01:22.62 +18:14:24.6 Longslit,1.65/210, seeing 0.75", 3600s, thin cirrus

12:55 UT --> flat, arcs, telluric

13:10 UT --> SDSS1138+0314 11:38:03.70 +03:14:58.0 Longslit, 1.93/200, seeing 0.6", 400s, thin cirrus, bright sky background on the guide probe towards the end. To be repeted

13:40 UT --> flat contaminated by sky line but no telluric, sky too bright

13:44 UT --> dome closed
I Attachment Action Size Date Who Comment
daq.20091211.jpgjpg daq.20091211.jpg manage 676 K 14 Dec 2009 - 23:16 JohnHill Plot of Telescope and Mirror Temperatures
pmcc.sx.thermal.20091211.jpgjpg pmcc.sx.thermal.20091211.jpg manage 503 K 14 Dec 2009 - 23:17 JohnHill Plot of SX Primary Mirror Temperatures
Topic revision: r5 - 14 Dec 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback